OpenStudy (anonymous):

3x + 6y = 18 and x + 2y = 6. What is the solution for the system? one solution no solution infinitely many solutions cannot be determined

OpenStudy (johnweldon1993):

put both into slope intercept form y =- 1/2x +3 and y = - 1/2x + 3 they're the same line...so there are an infinte number of solutions

OpenStudy (anonymous):

You really know how to go deeeep into.....your brain to find information dont you hmmm?(;good job